f) The price of petrol per litre has risen by 5%, 10%, 12%, 15% and 18% in each of five year successive years. Determine the geometric mean rise. (5 Marks)
1
Expert's answer
2019-03-11T12:28:25-0400

The geometric mean rise is used for estimating average growth of some values and can be calculated using the formula


"GM=\sqrt[n]{x_1\cdot x_2\cdot x_3\cdot ... \cdot x_{n}},"

which in our case (for the costs becoming 1.05, 1.1, ..., 1.18 times more expensive) gives


"GM=\sqrt[5]{1.05\cdot 1.10\cdot 1.12\cdot 1.15 \cdot 1.18}=1.119."

Therefore, every year the petrol costs 1.119 times more on the average than in the previous year (it adds 11.9% of its price for the past year).